Use the substitution method to find all solutions of the system of equations. (Round each answer to two decimal places.)
x^2 - y = 6
2x^2 + 3y = 16

Answer by checkley77(12844)

x^2 - y = 6 or y=x^2-6
2x^2 + 3y = 16 substitute (x^2-6) for y in the other equation & solve for x
2x^2+3(x^2-6)=16
2x^2+3x^2-18=16
5x^2=16+18
5x^2=34
x^2=34/5
x^2=6.8
x=sqrt6.8
x=2.61 ans.
y=(2.61)^2-6
y=6.8-6
y=.8 ans.
Proof:
2(2.61)^2+3*.8=16
2*6.8+2.4=16
13.6+2.4=16
16=16